ASTRONOMICAL SPECTROSCOPY: PAST SUCCESSES, FUTURE ENDEAVORS

E. HERBST, Department of Physics, The Ohio State University, Columbus, OH 43210; CARL A. GOTTLIEB, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138.

Molecular spectroscopy has played a decisive role in the study of both interstellar and circumstellar regions in astronomy. In this talk, we introduce the mini-symposium on ”New Trends in Astronomical Observations” by summarizing past successes, and looking forward to future endeavors.